# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

### v3

@@ Abstract
The past few years have produced a series of spectacular advances in the decoding of speech from brain activity. The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects. However, participants exhibit individual differences, such as anatomy, and datasets use varied scanners and task designs. As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets. In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning. This gap exists for all neural data, but especially for magnetoencephalography (MEG), where the scale of individual datasets has not yet caught up with other modalities. To address this, we develop a set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings. Experimental results with MEG show that representations learned with these objectives scale with data, generalise across subjects, datasets, and tasks, outperform using the raw input representation, and even surpass comparable self-supervised approaches. In addition, we set new benchmarks for two foundational speech decoding tasks. Collectively, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
Figure 1: Leveraging unlabelled data using pretext tasks for speech decoding. We pre-train a neural network using tasks that generate implicit labels from abundant unlabelled MEG neuroimaging data, permitting learning from large heterogeneous datasets. The tasks apply a randomly selected neuroscientifically relevant transformation $T$ to the data and the network predicts the transformation. We then train a linear probe on top of the pre-trained model, which remains frozen, with labelled data, achieving superior generalisation (cf. raw inputs) owing to the strength of the representation.
@@ 1 Introduction
In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases (Sutton, 2019). In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains (Jumper et al., 2021; Caron et al., 2021; OpenAI, 2023; Radford et al., 2023).
@@
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated. State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive (Moses et al., 2021; Willett et al., 2023) or non-invasive brain recordings (Tang et al., 2023), mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain. With few exceptions at present, e.g. Défossez et al. (2023), speech decoding models tend not to train on data from more than one subject. Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations, as has begun to happen in neighbouring sub-fields, such as  Jiang et al. (2024), holds the promise of training models on collective, internet-scale data.
@@
While neuroimaging modalities such as electroencephalography (EEG) are more abundant, MEG may be a better modality for decoding as it provides a richer signal (Lopes da Silva, 2013; Hall et al., 2014). Given the scarcity of speech-labelled MEG data and the relative abundance of other MEG data, self-supervised learning (SSL) appears promising as it is an avenue for domains where labels are rare or hard to obtain (Balestriero et al., 2023). But the scale of public MEG data, while large, is still not at the volume of breakthroughs in self-supervised image and natural language processing, let alone EEG. Thus, SSL methods for MEG need to be highly data-efficient. Pretext tasks are one such method in which domain-specific self-supervised tasks are used to pre-train a model on unlabelled data by generating implicit training labels through transformations of the input in order to help a downstream task. We develop a set of these tasks, informed by advances in neuroscience, for learning with unlabelled brain data (Figure 1) and design an architecture for processing continuous multi-sensor neuroimaging signals which we train using our pretext tasks. In order to scale existing non-invasive datasets, we provide a unified method that allows us to leverage data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners. We evaluate the representations learned with our approach on heard speech datasets acquired with non-invasive MEG, setting the baselines for speech detection and voicing classification on this data. The results not only demonstrate that scaling with unlabelled data works in speech decoding, but also shows that these representations can generalise across datasets, tasks, and even novel subjects for the first time. Our main contributions are:
@@
  • A set of domain-specific self-supervised pretext tasks for representation learning that can scale speech decoding over multiple subjects, multiple studies, and unlabelled data;
  • A data-efficient neural architecture for learning these self-supervised objectives and training downstream speech decoding from brain data; and
@@
  • A comprehensive experimental evaluation, using multiple times the volume of data in prior work, that verifies the above claims and additionally provides evidence for the existence of scaling laws when pre-training models with unlabelled MEG recordings.
@@ 2 Related Work
Prior work in speech decoding has focused almost entirely on supervised learning with decoding models that typically do not generalise across participants or experiments. This is true both in recent state-of-the-art invasive studies (Moses et al., 2021; Metzger et al., 2023; Willett et al., 2023; Chen et al., 2024) and non-invasive studies (Tang et al., 2023). These prior works have scaled up the experimental data collected within individual subjects, but are unable to leverage data from other subjects and experiments. Focusing on semantic rather than phonetic decoding, the method developed by Tang et al. (2023) is remarkable for showing an ability to generalise across labelled task data when listening to speech, imagining speech, or even watching videos. They do not, however, leverage unlabelled data and are unable to show generalisation between subjects.
@@
Specific studies into the limitations of generalising models between subjects show that while performance decreases on average when subjects are pooled, there are exceptions. Csaky et al. (2022) find that a subset of individuals perform better when evaluated with a group-level model than with individual models. Exploiting audio data in a multi-modal framework, Défossez et al. (2023) show that decoding performance improves for a segment identification task as data from multiple subjects listening to connected speech are aggregated. However, they do not demonstrate the ability to generalise to novel subjects and must retrain their model for new datasets. Moreover, although they repeat the result within two MEG and two EEG datasets, Défossez et al. (2023) do not show any improvements for pooling data across datasets. Their method is also unable to incorporate data without corresponding audio labels and so they do not combine data from studies with other kinds of labels either; cf. Wang & Ji (2022); Duan et al. (2023); Wang et al. (2023a). Unfortunately, the first two of these papers included a bug in their evaluation code. As such, their methods may perform no better than a baseline that provides pure noise inputs to the model (Jo et al., 2024).
[…]
@@
In the computer vision literature, there have been a plethora of methods that use self-supervised pretext tasks for representation learning (Agrawal et al., 2015; Doersch et al., 2015; Noroozi & Favaro, 2016; Larsson et al., 2016; Zhang et al., 2016; Gidaris et al., 2018). Until now, similar approaches have not translated to the brain decoding literature. However, prior work has used other methods to leverage unlabelled brain data. For example, Jiang et al. (2024) succeeded in cross-dataset and cross-task generalisation, using a transformer with tokenised brain signals and a masked token prediction objective. Although this work combined unlabelled datasets, their results studied simpler non-speech tasks with EEG. Wang et al. (2023b) used a similar approach, replacing tokens with contextualised embeddings of time-frequency input representations. Their impressive speech detection results were achieved with invasive neural recordings, which are comparatively rare and thus have much less potential to scale than non-invasive data. Perhaps the closest work to ours in terms of unlocking scaling with neural data is BIOT (Yang et al., 2023). This is a self-supervised architecture for encoding bio-signals that is similarly capable of training with different datasets, labels, and varied numbers of sensors. Like the previous works, the approach tokenises signals for a transformer architecture, but instead of a masked loss it uses a contrastive pre-training objective. While theoretically supporting MEG, Yang et al. (2023) evaluate BIOT on simple ECG/EEG tasks rather than address the comparatively complex challenge of speech decoding with MEG data.
@@ 3 Method
To encode continuous neuroimaging data, we introduce a neural architecture to embed heterogeneous brain signals. We leverage this architecture for self-supervised learning from unlabelled MEG data using a set of pretext tasks designed to generate generalisable brain representations for speech decoding. With this approach, we hope to replicate similar successes in computer vision (Gidaris et al., 2018; Chen et al., 2020).
@@ 3.1 Network Architecture
Our two-stage neural network architecture (Figure 2) uses pretext tasks in pre-training to learn a representation with unlabelled brain data. Then, the fine-tuning stage uses this representation to learn the downstream task by training with labelled data.
@@
We divide recordings into windows of length $w$ seconds or $t$ samples. At train time, each batch of windows is standardised such that each sensor has zero mean and unit variance. The network takes as input the standardised sample windows. To combine heterogeneous datasets, which have different numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$. Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs (Zeghidour et al., 2022; Défossez et al., 2022). Specifically, our convolutional encoder adapts the implementation of the SEANet architecture (Tagliasacchi et al., 2020) used in Défossez et al. (2022) which we describe here and as part of Figure 2. As these codecs typically operate on mono audio signals in $\mathbb{R}^{1\times t}$, while our signals are in $\mathbb{R}^{d_{\mathrm{shared}}\times t}$, we increase the convolutional channel dimension from $1$ to match $d_{\mathrm{shared}}$ while also inflating the channel dimension of subsequent convolutions. We refer to the output dimension of embeddings from this backbone as $d_{\mathrm{backbone}}$. Thus, the backbone takes as input a window in $\mathbb{R}^{S\times t}$, and encodes this into $\tau$ embeddings (where $\tau<t$), each of dimension $d_{\mathrm{backbone}}$ (i.e. an $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ output).
Just as speakers have different voices, neural responses between subjects have different characteristics. Consequently, individual variation leads to models that do not generalise well across subjects (Csaky et al., 2022). In the speech literature, models include speaker conditioning to account for these differences (Gibiansky et al., 2017). We take a similar approach by introducing subject conditioning. Zeghidour et al. (2022) find that conditioning is equally effective at the encoder bottleneck as in other stages of the model. Hence, we place ours at the cortex encoder bottleneck for simplicity. We use feature-wise linear modulation (FiLM) (Perez et al., 2018) as our conditioning method.
Following the advice of Balestriero et al. (2023, Section 3.2), we use a two-layer feedforward projector to alleviate misalignment between our pretext and downstream tasks in the representation. After the projector, linear classifiers make predictions for each of the pretext tasks. When fine-tuning, we train a linear decoder, for a downstream task, on top of the pre-trained representation, which remains frozen. Thus, we backpropagate only through the classifier. A trainable dataset-specific linear layer can be introduced for a novel dataset.
@@ 3.2 Pretext Tasks
Our pretext tasks are unsupervised feature learning tasks that aim to learn generalisable speech decoding features. Since different datasets use varied numbers of sensors, we construct these tasks with labels that are agnostic to the number of sensors in the signal.
Band prediction. In the literature, neural responses can be segmented into functional frequency bands (Giraud & Poeppel, 2012; Piai et al., 2014; Mai et al., 2016). Delta ($\delta$) waves (0.1–4 Hz) are commonly associated with the rhythmic structure of heard speech (Luo et al., 2010), Theta ($\theta$) waves (4–8 Hz) reliably track (Luo & Poeppel, 2007) and phase-lock to the amplitude envelope of heard sentences (Peelle et al., 2012), Alpha ($\alpha$) waves (8–12 Hz) relate to attentional processes and the inhibition of irrelevant information, helping to focus on relevant speech signals (Strauß et al., 2015), Beta ($\beta$) waves (12–30Hz) are implicated in top-down predictive coding (Bressler & Richter, 2015) which affects lexical processing (Weiss & Mueller, 2012), Gamma ($\gamma$) waves (30–70 Hz) occur with higher cognitive functions (e.g. memory, learning, reasoning, and planning) (Fries, 2009; Buzsáki & Wang, 2012), and High Gamma ($\gamma^{\mathrm{high}}$) waves ($>$70 Hz) have been linked specifically to speech detection (Hamilton et al., 2018) and phonemic feature classification in the STG (Mesgarani et al., 2014) as well as phonemic feature classification in the ventral sensorimotor cortex (vSMC) (Cheung et al., 2016). As High Gamma is a relatively wide band, we have split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz).
To learn representations that can distinguish between these, our band prediction task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding linear predictor $f_{\mathrm{band}}$, requiring the network to predict the frequency band that was rejected. This yields the loss
@@
Phase shift prediction. Phase coupling between networks of neuron populations is necessary for coordinating brain activity (Fries, 2005; Vidaurre et al., 2018). Thus, since phase often synchronises between communicating brain areas, phase coupling between spatially distant sensors is likely to be a useful feature. Supporting this insight, recent work (Jiang et al., 2024) also finds phase to be an essential component of the signal.
@@
These pretext tasks capture complementary time- and frequency-domain properties of the signal. Hence, during pre-training, we combine them, creating an augmented version of the input for every pretext task by applying the matching transformation. We feed the augmented inputs through the network backbone and apply the corresponding classifier to predict the transformation, summing the weighted losses such that our final pre-training loss is given by

### v4

@@ Abstract
The past few years have seen remarkable progress in the decoding of speech from brain activity, primarily driven by large single-subject datasets. However, due to individual variation, such as anatomy, and differences in task design and scanning hardware, leveraging data across subjects and datasets remains challenging. In turn, the field has not benefited from the growing number of open neural data repositories to exploit large-scale deep learning. To address this, we develop neuroscience-informed self-supervised objectives, together with an architecture, for learning from heterogeneous brain recordings. Scaling to nearly 400 hours of MEG data and 900 subjects, our approach shows generalisation across participants, datasets, tasks, and even to novel subjects. It achieves improvements of 15-27% over state-of-the-art models and matches surgical decoding performance with non-invasive data. These advances unlock the potential for scaling speech decoding models beyond the current frontier.
Machine Learning, ICML
Figure 1: Leveraging unlabelled data using pretext tasks for speech decoding. We pre-train a neural network using tasks that generate implicit labels from abundant unlabelled MEG neuroimaging data, permitting learning from large heterogeneous datasets. The tasks apply a randomly selected neuroscientifically relevant transformation $T$ to the data and the network predicts the transformation. We then train a linear probe on top of the pre-trained model, which remains frozen, with labelled data, achieving superior generalisation owing to the strength of the representation.
@@ 1 Introduction
In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases (Sutton, 2019). The ability of deep learning to learn from ever-larger datasets has enabled seemingly arbitrary scaling with computation, leading to astounding advances across a diverse set of domains (Jumper et al., 2021; Caron et al., 2021; OpenAI, 2023; Radford et al., 2023).
@@
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated. Recent work towards brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive (Moses et al., 2021; Willett et al., 2023) or non-invasive brain recordings (Tang et al., 2023), mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain, especially in magnetoencephalography (MEG) data. Current speech decoding models rarely train on multiple subjects, combine datasets, or utilise data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations holds the promise of training models on collective, internet-scale data.
@@
Decoding methods for MEG need to be highly data-efficient. While electroencephalography (EEG) data are abundant, MEG provides richer signals for decoding (Lopes da Silva, 2013; Hall et al., 2014) but are comparatively rare. Given the scarcity of speech-labelled MEG and the larger proportion of other MEG data, self-supervised learning (SSL) appears promising as it is an avenue for domains where labels are rare or hard to obtain (Balestriero et al., 2023). To data-efficiently learn from unlabelled MEG, we propose pretext training with neuroscience-informed input transformations that benefit downstream tasks. We use this for learning from unlabelled brain data (Figure 1) through an architecture for processing continuous multi-sensor neuroimaging signals. Our method provides a unified approach that enables leveraging data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners. We evaluate representations learned with our approach on heard speech datasets acquired with MEG, setting the baselines for speech detection and voicing classification on this data.
Our main contributions are:
@@
  • A domain-specific self-supervision method and a neural architecture for representation learning from MEG that unlock scaling speech decoding over multiple subjects, multiple studies, and unlabelled data;
  • Achieving 15-27% gains over state-of-the-art self-supervised models, matching surgical self-supervised decoding non-invasively, and showing novel subject generalisation for the first time in MEG; and
@@
  • Demonstrating evidence for scaling laws arising from pre-training with unlabelled MEG recordings using multiple times the volume of data in prior work.
@@ 2 Related Work
Prior work in speech decoding has focused almost entirely on supervised learning with decoding models that typically do not generalise across participants or experiments. This is true both in recent state-of-the-art invasive studies (Moses et al., 2021; Metzger et al., 2023; Willett et al., 2023; Chen et al., 2024a) and non-invasive studies (Tang et al., 2023). These prior works have scaled up the experimental data collected within individual subjects, but are unable to leverage data from other subjects and experiments. Nevertheless, the method developed by Tang et al. (2023) is remarkable for showing an ability to generalise across labelled task data. They do not, however, use unlabelled data or show cross-subject generalisation.
@@
Specific studies into the limitations of generalising models between subjects show that while performance decreases on average when subjects are pooled, there are exceptions (e.g. Anumanchipalli et al. (2019) and Makin et al. (2019) in surgical settings and Csaky et al. (2022) non-invasively). Défossez et al. (2023) show cross-subject generalisation for a segment identification task from participants listening to connected speech. However, they do not demonstrate generalisation to novel subjects and retrain their model for new datasets rather than being able to generalise across datasets or pool them. Their method is also unable to incorporate data without corresponding audio labels and so does not scale with other kinds of tasks.
In general, speech decoding has centred on different kinds of speech: listening, imagining, speaking out loud, and, for paralysed patients, attempting to speak aloud. We focus on listening because it is easier to decode than imagined speech (e.g. Martin et al. (2014)). There is also evidence of functional overlap between listening and imagined speech representations in the brain (Wandelt et al., 2024), though the question of overlap has been contested (Langland-Hassan & Vicente, 2018). While some work on decoding text directly from heard speech tasks with MEG and EEG exist, it is unclear whether these methods perform any better than a baseline that provides pure noise inputs to the model (Jo et al., 2024). Non-invasive speech decoding remains a highly challenging and unsolved domain.
@@
Self-supervised pretext tasks have been successful in computer vision (Agrawal et al., 2015; Doersch et al., 2015; Noroozi & Favaro, 2016; Larsson et al., 2016; Zhang et al., 2016; Gidaris et al., 2018) but rarely applied to brain decoding. There are, however, methods that leverage unlabelled brain data in other ways (Banville et al., 2019; Kostas et al., 2021; Le & Shlizerman, 2022; Zhang et al., 2023; Yi et al., 2023; Cai et al., 2023; Ye et al., 2023; Yuan et al., 2024; Chen et al., 2024b). Unfortunately, most of this literature is unable to scale and harmonize heterogeneous non-invasive data. The most notable of these works include Wang et al. (2023), who learn contextualised embeddings of time-frequency input representations through masked spectrogram in-filling. Their impressive speech detection results were achieved with invasive neural recordings, which are comparatively rare and thus have less potential to scale than non-invasive data. Another, BIOT (Yang et al., 2023), learns from generic heterogeneous bio-signals with a contrastive pre-training objective, rather than masking, and applies it to ECG/EEG data. Notably, none of these methods optimise their objectives for speech decoding or focus on MEG.
@@ 3 Method
We introduce a neural architecture to embed heterogeneous brain signals. Then, we leverage this architecture for self-supervised learning from unlabelled MEG data using a set of pretext tasks designed to generate generalisable brain representations for speech decoding.
@@ 3.1 Network Architecture
Our neural network architecture has two stages (Figure 2): pre-training with pretext tasks on unlabelled data, and training a linear probe with labelled data for downstream tasks.
@@
We divide recordings into windows of length $w$ seconds or $t$ samples. At train time, each batch of windows is standardised such that each sensor has zero mean and unit variance. The network takes as input the standardised sample windows. To combine heterogeneous datasets, which have varying numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$. Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs (Zeghidour et al., 2022; Défossez et al., 2022). Specifically, our convolutional encoder adapts the SEANet architecture (Tagliasacchi et al., 2020) used in Défossez et al. (2022) which we describe here and as part of Figure 2. As these codecs typically operate on mono audio signals in $\mathbb{R}^{1\times t}$, while our signals are in $\mathbb{R}^{d_{\mathrm{shared}}\times t}$, we increase the convolutional channel dimension from $1$ to match $d_{\mathrm{shared}}$ while also inflating the channel dimension of subsequent convolutions. We refer to the output dimension of embeddings from this backbone as $d_{\mathrm{backbone}}$. Thus, the backbone takes as input a window in $\mathbb{R}^{S\times t}$, and encodes this into $\tau$ embeddings, each of dimension $d_{\mathrm{backbone}}$ (i.e. an $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ output).
In the speech recognition literature, models include speaker conditioning to account for vocal and prosodic differences (Gibiansky et al., 2017). Just as speakers have different voices, neural responses between subjects have different characteristics. Consequently, individual variation leads to models that do not generalise well across subjects (Csaky et al., 2022). We address this with a similar approach to the speech literature by introducing subject conditioning using feature-wise linear modulation (FiLM) (Perez et al., 2018). As Zeghidour et al. (2022) find that conditioning is as equally effective at the encoder bottleneck as in other stages of the model, we also condition at the bottleneck.
Following Balestriero et al. (2023, Section 3.2), we use a two-layer projector to alleviate misalignment between our pretext and downstream tasks in the representation. After the projector, linear classifiers make predictions for each of the pretext tasks. When fine-tuning, we train a linear decoder, for a downstream task, on top of the pre-trained representation, which remains frozen. Thus, we backpropagate only through the classifier. A trainable dataset-specific linear layer can be introduced for a novel dataset.
@@ 3.2 Pretext Tasks
To use our architecture for pre-training, we construct pretext objectives for unsupervised learning of generalisable speech decoding features. These objectives are inherently agnostic to the sensor count because they operate on properties that are independent of the specific sensor arrangement. This key design choice enables the tasks to work seamlessly across datasets with varying numbers of sensors—a critical requirement for combining heterogeneous brain data.
Band Hz Association
Delta ($\delta$) .1-4 Rhythmic structure of heard speech (Luo et al., 2010)
Theta ($\theta$) 4-8 Tracking (Luo & Poeppel, 2007) and phase-locking to the amplitude envelope of heard sentences (Peelle et al., 2012)
Alpha ($\alpha$) 8-12 Attentional processes and the inhibition of irrelevant information (Strauß et al., 2015)
Beta ($\beta$) 12-30 Top-down predictive coding (Bressler & Richter, 2015) which affects lexical processing (Weiss & Mueller, 2012)
Gamma ($\gamma$) 30-70 Higher cognitive functions (e.g. memory, learning, reasoning, and planning) (Fries, 2009; Buzsáki & Wang, 2012)
High 70+ Speech detection (Hamilton et al., 2018) and phonemic feature classification in the STG (Mesgarani et al., 2014) and the ventral sensorimotor cortex (vSMC) (Cheung et al., 2016)
Gamma
($\gamma^{\mathrm{high}}$)
Table 1: Functional frequency bands in brain activity.
Band prediction. In the literature, neural responses can be segmented into functional frequency bands (Giraud & Poeppel, 2012; Piai et al., 2014; Mai et al., 2016) (Table 1). Sensitivity to these frequencies would bring about functional separability in the representation space. Thus, to learn such representations, we train the network to classify rejected bands. As High Gamma is a relatively wide band we split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz). Our task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding linear predictor $f_{\mathrm{band}}$, requiring the network to classify which frequency band $\omega$ was rejected. This yields the loss
@@
Phase shift prediction. Phase coupling between networks of neuron populations is necessary for coordinating brain activity (Fries, 2005; Vidaurre et al., 2018) and so phase coupling between spatially distant sensors is likely to be a useful feature. Supporting this insight, recent work (Jiang et al., 2024) also finds phase to be an essential component of the signal.
@@
Combined tasks. These pretext tasks capture complementary time- and frequency-domain properties of the signal. Hence, during pre-training, we combine them, creating an augmented version of the input for every pretext task by applying the matching transformation. We feed the augmented inputs through the network backbone and apply the corresponding classifier to predict the transformation, summing the weighted losses such that our final pre-training loss is